OpenStudy (anonymous):

evaluate to make a rational number log5(3√625)−9^(3/2) 5- is the base..

OpenStudy (anonymous):

the last part is easiest \[9^{\frac{3}{2}}=\sqrt{9}^3=3^3=27\]

OpenStudy (anonymous):

the first part you can use \[\frac{1}{3}\log_5(625)=\frac{1}{3}\log_5(5^4)=\frac{4}{3}\]